OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1 A. -69x + 10y = 51 B. 69x + 10y = 51 C. 69x + 10y = -51 D. -69x + 10y = -51

OpenStudy (anonymous):

Is the original equation y = -6.9x + 5.1 or y = -69x + 51

OpenStudy (anonymous):

y = -6.9x + 5.1

OpenStudy (anonymous):

Ok try each one till you fin one that fits: A. wrong -69x + 10y = 51 10y = 69x + 51 y = 6.9x + 5.1 B. 69x + 10y = 51 10y = -69x + 51 y = -6.9x + 51 So whats your answer?

OpenStudy (anonymous):

How about multiply both sides for 10: 10y = ... @alyssababy7 Can you continue?

OpenStudy (anonymous):

That's easier^

OpenStudy (anonymous):

@jazy sighs, still no hope for @alyssababy7

OpenStudy (anonymous):

i'm sorry i lose connection a lot at my house.

OpenStudy (anonymous):

multiply both sides for 10: 10y = .... Can you continue?